OpenStudy (anonymous):

Luke raises rabbits and raccoons to sell for meat. The price of raccoon meat is three times the price of rabbit meat. One day luke sold 160 pounds of meat, $72 worth of each type. What is the price per pound of each type of meat?

OpenStudy (dumbcow):

Let A be price of raccoon meat, B the price of rabbit meat, x is pounds of raccoon meat sold A = 3B A*x = 72 B*(160-x) = 72

OpenStudy (dumbcow):

Use substitution 3B*x = 72 Bx = 24 ---------------- 160B -Bx = 72 160B - 24 = 72 160B = 96 B = 96/160 = 6/10 = 0.6 ------------------ A = 3*0.6 = 1.8